A deuteron and alpha particle are placed in an electric feild .The force acting on them are F1 and F2 and acceleration a1 and a2 respectively . Then

A deuteron and  alpha  particle are placed in an electric feild .The force acting on them are F1 and F2  and acceleration  a1 and a2 respectively . Then
 
 

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Hiii 
For Deutron q = 1
M1 = 2
For alpha q= 2
M2 = 4
F1 for deutron = qe 
F2 for alpha = qe = 2*e = 2e
a1 for Deutron = F1 /M1
=> e/2
a2 for alpha = 2e /4 = e/2 
a1 = a2